OpenStudy (anonymous):

1) Explain what a 30-year, 5/3 ARM at 3.25% with a 2/9 cap means

OpenStudy (anonymous):

can someone please help me!

OpenStudy (anonymous):

@shifuyanli

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

is mortgage stuff

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

I have 6 questions, could you help me with it?

OpenStudy (anonymous):

I certainly can't to be honest but I can tag some people to help you :)

OpenStudy (anonymous):

yes please!!! I need to do it ASAP

OpenStudy (anonymous):

@iGreen @AnswerMyQuestions @cecormier @YanaSidlinskiy @Jaynator495

OpenStudy (anonymous):

@GreenCat

OpenStudy (anonymous):

Please help me please!!!

OpenStudy (greencat):

@amistre64

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

1) Explain what a 30-year, 5/3 ARM at 3.25% with a 2/9 cap means

OpenStudy (anonymous):

amistre is here

OpenStudy (amistre64):

tell me your thoughts on it .. ill correct as needed

OpenStudy (anonymous):

I just really dont know, not even where to start my thoughts

OpenStudy (amistre64):

well, it a loan .. how long is it?

OpenStudy (anonymous):

30 years ?

OpenStudy (amistre64):

yep, now is it a fixed rate or an adjustable rate?

OpenStudy (anonymous):

adjustable

OpenStudy (amistre64):

so we know its a 30 year loan that has a rate that can change ... we need to know what the terms of that change is. 5/3 tells us about how often and when to make adjustments. can you tell me what you think it would mean?

OpenStudy (anonymous):

im not sure :/

OpenStudy (amistre64):

if youve read your material you should have some inkling of a notion ... if not, then just make something up that sounds reasonable ... i cant do the thinking for you, but i can correct it.

OpenStudy (amistre64):

we have a 30 year loan, that changes its rate; 5/3 tells us when and how often to change it. come up with a notion that you think it may be ... its fine to be wrong

OpenStudy (anonymous):

would it mean 5 time every 3 years?

OpenStudy (amistre64):

thats at least a thought :) the rate stays the same for the first 5 years of the loan afterwards, every 3 years we are allowed to make an extra adjustment 5 initial years -------------------- 3 years for adjustments now tell me what i said :)

OpenStudy (anonymous):

we have a loan of 30 years that has a rate that can change. It stays the same the first 5 years, and then it can adjust every 3 years

OpenStudy (amistre64):

..every 3 years afterwards. good now we know how often it changes, we need to know HOW it changes what is the initial interest rate?

OpenStudy (anonymous):

3.25%?

OpenStudy (amistre64):

good, now we have to adjust it based on the terms 2/9 cap since this is telling us about the interest rate change, what are your thoughts as to its meaning?

OpenStudy (anonymous):

it stays the same for the first 2 years and then it changes every 9 years?

OpenStudy (amistre64):

no, but im glad your applying previous thoughts. it is similar but: when we change the rate: it can only be raised up to 2% each time and all the adjustments made, cannot exceed 9% in total. lets say im giving you candy: every hour for the first 5 hours, i give you 3.25 pieces of candy then i add 2 more pieces, for the next 3 years, you get 5.25 pieces of candy then i add 2 more pieces, for the next 3 years, you get 7.25 pieces of candy then i add 2 more pieces, for the next 3 years, you get 10.25 pieces of candy i can only give you 9 extra pieces all together, ive given you 8 so far soo for the rest of your life i give you 11.25 pieces of candy each year.

OpenStudy (amistre64):

i chnged hours to years lol .... keep it all hours, or keep it all years .... concepts the same tho

OpenStudy (anonymous):

ohhhok

OpenStudy (anonymous):

so then will all these concepts its what those terms means? 1) Explain what a 30-year, 5/3 ARM at 3.25% with a 2/9 cap means cus this is the question^^^

OpenStudy (amistre64):

yep for 30 years, we pay back a loan that has an adjustable rate starting at 3.25% the rate stays that same for the first 5 years, and then can only change every 3 years afterwards the rate changes by at most by 2%, and the combined total adjustments cannot exceed 9% extra, so we are capped at 3.25+9 = 12.25 for the remainder of the loan.

OpenStudy (anonymous):

yay!!!

OpenStudy (anonymous):

I have 5 more questions but i dont want to take your time are you busy?

OpenStudy (amistre64):

>(B)< have a butterscotch :) i can prolly help with one more

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

can it be 2 more? <3

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

maybe, lets try one ore and see how my sanity holds up lol

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

2) Explain the process for finding the remaining balance of a condo costing $200,000 and using the terms from question #1. *Please don’t discuss the actual calculation, but the steps or process you go through to find the remaining balance and what that amount represents.

OpenStudy (amistre64):

the same way we would determine the loan balance normally. do you know a formula that determines the amount of a loan owed?

OpenStudy (amistre64):

this would be an essay if we have to explain how a loan calculation functions as a matter of time ...

OpenStudy (anonymous):

isnt it the full price minus what was already paid?

OpenStudy (anonymous):

I only have to write the steps right?

OpenStudy (amistre64):

regretabbly no the payment is a combination of principal and interest; all the while the balance is compounding up and up if the payments dont at least cover the interest that has collected, then we can never pay of the loan.

OpenStudy (amistre64):

if we had a formula predefined for us, we could use it at every point in time that we need to make an adjustement.

OpenStudy (anonymous):

so you dont know how i could answer this?

OpenStudy (amistre64):

i have a formula that i developed .. but im not sure if its the same one that you would use.

OpenStudy (amistre64):

i know how i would answer it, but im not taking the class so i dont know how YOU should answer it.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

we can use your formula

OpenStudy (amistre64):

if we had a formula Bn that calculates the amount of loan to repay, then we would know how to adjust the terms of the loan. after 5 years: B5 is a new loan, offered at the adjusted rate. 3 years afterwards: B8 is a new loan, offered at the adjusted rate. 3 years afterwards: B11 is a new loan, offered at the adjusted rate. 3 years afterwards: B14 is a new loan, offered at the adjusted rate. 3 years afterwards: B17 is a new loan, offered at the adjusted rate. from B17 to B30, the rate doesnt change

OpenStudy (amistre64):

i think the sites crashing again ... or my sanity is failing :)

OpenStudy (anonymous):

yeah, its bead right now

OpenStudy (anonymous):

bad****

OpenStudy (amistre64):

not without explaining it, but i think thats not quite what the question is asking for. the balance of a loan at any given payment period can be determined by how much interest has accumulated, and how many payments have been made. B0 = B0 , we take out a loan for some Balance B, it takes on interest of a compounding rate of 'k' B1 = B0(k) - P , we make a payment, after the interest add up ----------------------------- B2 = B1(k) - P , we make a payment, after the interest add up but ... we know the value of B1, use it B2 = (B0(k) - P)(k) - P B2 = B0(k)^2 - P(k) - P ------------------------------- B3 = B2(k) - P , we make a payment, after the interest add up but ... we know the value of B2, use it B3 = (B0(k)^2 - P(k) - P)k - P B3 = B0(k)^3 - P(k)^2 - P(k) - P ------------------------------ this carries on for the life of the loan and it forms a pattern that we can turn into a formula Bn = Bk^n -P(k^n-1)/(k-1) but im pretty sure this is NOT what your expected to explain

OpenStudy (anonymous):

oh :/

OpenStudy (amistre64):

if we know the formula for the balance of a loan to begin with, we just say we use it ... if we had a formula Bn that calculates the amount of loan to repay, then we would know how to adjust the terms of the loan. after 5 years: B5 is a new loan, offered at the adjusted rate. 3 years afterwards: B8 is a new loan, offered at the adjusted rate. 3 years afterwards: B11 is a new loan, offered at the adjusted rate. 3 years afterwards: B14 is a new loan, offered at the adjusted rate. 3 years afterwards: B17 is a new loan, offered at the adjusted rate. for the remining year of the loan stays the same

OpenStudy (amistre64):

**remaining years

OpenStudy (anonymous):

3) Explain the process for finding the balloon payment for a house costing $190,000. The mortgage is a 20/5 Balloon at 4.05%. this is the other question

OpenStudy (amistre64):

how do you define a balloon payment? what does this type of loan represent to you?

OpenStudy (anonymous):

ima start another quesito and tag you ok?

OpenStudy (amistre64):

20/5 means the loan is figured as if it was a 20 year fixed rate loan the loan is paid for 5 years, then the loan is paid off with a balloon payment. the balloon payment is the value of the amount of the loan after 5 years.

OpenStudy (amistre64):

ill be back later, the sites not acting right on my end at least. tag me and ill be able to find it later ... if this is a timed event, i wish you luck tho :)